OpenStudy (anonymous):

7/1 + 1/9 = 1/9 + 1/6 = 6/9 + 1/3 = 8/10 - 1/20 = 6/7 - 4/21 = 6/11 - 1/5 =

OpenStudy (zepp):

Put them under the same denominator 7/1 + 1/9 would be 7*9/1*9 + 1/9 = 63/9 + 1/9 = 64/9